How much heat is needed to change 12.0 g of mercury at 20∘C into mercury vapor at the boiling point?

Express your answer with the appropriate units.
Answer #1

Solution)

We know, specific heat Mercury is 140 J/kgC
Mercury 11.3 kJ/kg (melts at –38.9ºC)
Heat of vaporization 295 kJ/kg
Boiling point 629.88 K, 356.73 °C

Energy needede to heat the Hg from 20 to 356.73ºC
E1 = 140 J/kgC x 0.012 kg x (356.73–20)C

E1= 565.70 J

Now, energy to boil the Hg
E2 = 295000 J/kg x 0.012 kg

E2= 2.45*10^7 J

Total Energy= E1+ E2

=565.70 J + 2.45*10^7 J = 2.45 *10^7 J (Ans)

=========

Good luck!:)

> How tf did you get 2.45*10^7 from 295000 x .012

Ryan Jackson Wed, Feb 16, 2022 10:03 PM